OpenStudy (anonymous):

find the θ values for the points of intersection for the curves r=3+sinθ and r=3-sinθ

OpenStudy (whpalmer4):

\[3 + \sin \theta = 3 - \sin \theta\] Solve for \(\theta\) and write an expression that reflects all of the solutions.

OpenStudy (whpalmer4):

Why did you close this and open another?